An NCC parade is going at a uniform speed of 6 km/h through a place under a berry tree on which a bird is sitting at a height of 12.1 m. At a particular instant the bird drops a berry. Which cadet (give the distance from the tree at the instant) will receive the berry on his uniform?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Let the true of fall be ‘t’ initial velocity u = 0 Acceleration a = 9.8 m/s2 Distance S = 12/1 m ∴ S = ut + ½ at2 ⇒ 12.1 = 0 + 1/2 (9.8) × t2 ⇒ t2 = 12.1/4.9 = 2.46 ⇒ t = 1.57 sec For cadet velocity = 6 km/hr = 1.66 m/sec Distance = vt = 1.57 × 1.66 = 2.6 m. The cadet, 2.6 m away from tree will receive the berry on his uniform.